OpenStudy (anonymous):

Given the function f(x) = 4(x+3) – 5, solve for the inverse function when x = 3.

OpenStudy (anonymous):

@johnweldon1993 this is my last question, if you don't mind!

OpenStudy (johnweldon1993):

Alright, so here...we need the inverse...and then plug in that value x = 3 have you done inverses before?

OpenStudy (anonymous):

Yes but they are VERY new to me. This is my first year of Algebra 2.

OpenStudy (johnweldon1993):

Ah, well welcome, it only gets harder from here >.< lol no, just kidding :P So For inverses...all we need to do...is switch the positions of the 'x' and the 'y'...and then solve for the 'y' again

OpenStudy (johnweldon1993):

So here we have \[\large y = 4(x + 3) - 5\]

OpenStudy (johnweldon1993):

Switch the 'x' and the 'y' \[\large x = 4(y + 3) - 5\]

OpenStudy (johnweldon1993):

and use algebra to solve for 'y' again... how would we do that?

OpenStudy (anonymous):

Combine terms (not sure if that's the proper term) to get 4y+7?

OpenStudy (johnweldon1993):

Indeed :) so we have \[\large x = 4y + 7\] so solving that for 'y' we get?

OpenStudy (anonymous):

x-4y=7?

OpenStudy (johnweldon1993):

Not quite :) so what the whole idea is...is to isolate the 'y' because eventually we want y = ...something right? So where we have \[\large x = 4y + 7\] We want y by itself....so we need to move everything else to the other side of the equation If we subtract 7 from both sides that will isolate the 4y...so that's a start! \[\large x - 7 = 4y\] Now...if we want to solve for y (not 4y) what do we do?

OpenStudy (anonymous):

Hm. I'm not sure.

OpenStudy (johnweldon1993):

Well... \[\large 4y = x - 7\] Lets see...we have a 4 being MULTIPLIED to the 'y'....so to cancel that...we do what is opposite of multiplication...which is division Since 4 is being multiplied to the y...if we divide that by 4...we will have \[\large \frac{4y}{4} = y\] which is what we want right? we wanted to just have 'y' So that's what we need to do...we need to divide everything by 4 \[\large \frac{x - 7 = 4y}{4}\] \[\large \frac{x - 7}{4} = y\]

OpenStudy (anonymous):

Ohh right!

OpenStudy (johnweldon1993):

Now, that is the whole process you will ALWAYS use for inverses...so does that make sense? any parts cloudy?

OpenStudy (anonymous):

The given answers are single numbers, though.

OpenStudy (johnweldon1993):

*We aren't done yet ^_^

OpenStudy (anonymous):

OH alright!

OpenStudy (johnweldon1993):

Remember we still have the "when x = 3" part to deal with :)

OpenStudy (johnweldon1993):

SOOOOOO since we know that \[\large y = \frac{x - 7}{4}\] "when x = 3" what does y equal?

OpenStudy (anonymous):

-1?

OpenStudy (johnweldon1993):

Indeed! :)

OpenStudy (anonymous):

Thank you so very much!

OpenStudy (johnweldon1993):

Of course, Anytime! :)